OpenStudy (anonymous):

Solve for g. one-twelfth equals g minus five-sixths g equals negative eleven-twelfths g equals negative three-fourths g equals three-fourths g equals eleven-twelfths

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

1/12 = g - 5/6 (add 5/6 to each side) 1/12 + 5/6 = g (common denominator is 12) 1/12 + 10/12 = g 11/12 = g or if you do not want to mess with fractions... 1/12 = g - 5/6 (multiply the entire equation by LCD which is 12) 1 = 12g - 10 (add 10 to both sides) 10 + 1 = 12g 11 = 12g (now divide by 12) 11/12 = g